OpenStudy (anonymous):

What number(s) should be excluded as values of x in the polynomial? x+2 over x^2-4

OpenStudy (anonymous):

this is not a polynomial, its a rational expression. Since dividing by 0 is undefined, you need to find the values of x that make this expression undefined, which will happen when x^2-4=0

jhonyy9 (jhonyy9):

x+2 ---------- = 1/(x-2) (x+2)(x-2)